# Position: The Right to AI

### v1

@@ 2.4 Grassroots Engagement
Previous attempts to incorporate participation in AI governance include user feedback loops in recommender systems, collaborative training data annotation, and community reviews of AI outputs (Gerdes, 2022; Zaidan & Ibrahim, 2024; Huang et al., 2024a). Innovative proposals like “jury-based” or “constitutional” approaches also engage diverse groups in AI ethics and policy deliberations (Gordon et al., 2022; Bai et al., 2022b; Sorensen et al., 2024). However, these methods are nascent and face scalability, political, market, resource, and institutional challenges (Saheb & Saheb, 2024; Zaidan & Ibrahim, 2024; Huang et al., 2024a; Stammbach et al., 2024).
@@ 3 Arguments for AI as Societal Infrastructure
A central premise of the Right to AI is that AI increasingly functions as societal infrastructure, comparable to utilities or educational systems. Viewing AI as societal infrastructure aligns with established frameworks of public goods, commons governance, and socio-technical systems (Ostrom, 1990; Graham & Marvin, 2001). Infrastructure commonly exhibits three properties: (i) broad societal impact, (ii) an essential role in daily life, and (iii) a requirement for collective management (North, 1990; Davern et al., 2017). AI, particularly foundation models shaping decisions having to do with employment, credit scoring, and public discourse, for example, arguably meets these criteria (Eubanks, 2018; Gillespie, 2018; Angwin et al., 2022).
@@
Technologies that mediate access to financial systems, public services, and employment increasingly function as core societal infrastructure (North, 1990; Gary Bridge, 2011). AI-driven decision-making in credit approval, job screening, and social welfare administration underscores its role in structuring life chances (Eubanks, 2018; Benjamin, 2019). The opacity of these systems necessitates governance mechanisms akin to those regulating financial and legal infrastructures (Ulnicane, 2024).
@@
AI-based systems shape social interactions, political communication, and institutional trust (Gillespie, 2018; Kitchin, 2023). Like other infrastructures, AI is not neutral; it embeds political, economic, and cultural assumptions that shape its societal consequences (Graham & Marvin, 2001; Montalto et al., 2019; Angwin et al., 2022). Without participatory oversight, AI risks reinforcing inequities rather than serving as a mechanism for collective well-being.
@@ 4.2 Social Justice and Pluralism
Machine learning models typically generalize from large datasets, which may fail to capture minority values or nuanced cultural norms (Goodfellow et al., 2016; Gebru et al., 2021; Dastin, 2022). As a result, marginalized voices risk erasure or misrepresentation (Raz, 1999; Beer, 2016; Bondi et al., 2021; Zhang et al., 2025). The Right to AI entails inclusive governance structures that protect pluralism by ensuring that multiple moral and cultural frameworks inform system design (Lefebvre, 1968; Costanza-Chock, 2020). This pluralistic perspective challenges any hegemonic assumption that there is a single “correct” data-driven solution (Fraser, 1995; Ostrom, 1996; Kitchin, 2023; Sorensen et al., 2024).
@@ 5.1 Consumer-Based (Minimal Right to AI)
5.2 Private Organization-Led (Proprietary Technocracy)
@@
5.3 Government-Controlled (Bureaucratic Sovereignty)
@@
Figure 2: Progression in stakeholder power from minimal engagement (Extractive Consumerism) to robust self-governance (Communal Sovereignty). This categorization helps assess current initiatives and guide transitions toward more participatory models.
@@ 5.4 Citizen-Controlled (Maximal Right to AI)
Figure 2 illustrates how agency, transparency, inclusivity, and governance structures vary across these tiers. This progression also highlights the transition from extractive consumerism to communal sovereignty, guiding evaluations of existing approaches and helping chart paths toward more participatory paradigms.
@@ 6 Lessons from Participatory Practices
Participatory AI initiatives in education, healthcare, urban planning, and software development often involve stakeholders with direct interests in AI-driven decisions (Jacobs, 1961; Lee et al., 2019; Zicari et al., 2021; Zhang & Aslan, 2021; Sieber et al., 2024b). For example, Co-Design of Trustworthy AI in Healthcare (Zicari et al., 2021) included patients, clinicians, and ethicists to expose biases in diagnostic tools, leading to enhanced accountability despite resource and expertise challenges. Urban planning efforts, such as MID-Space (Nayak et al., 2024), relied on iterative community annotation and mediation to address conflicting priorities (Sloane et al., 2022). Similarly, WeBuildAI (Lee et al., 2019) employed workshops to refine algorithmic governance norms, decentralizing certain design choices. These examples highlight both the promise of stakeholder inclusion and the structural, institutional, and practical barriers that may limit its impact.
@@ 8 Alternative Views
Some scholars and practitioners question whether broad-based participatory approaches to AI are feasible or desirable. A market-led perspective asserts that competition and consumer choice will naturally drive responsible AI (Dignam, 2020; de Marcellis-Warin et al., 2022; Hadfield & Clark, 2023; Judge et al., 2024), though such models can overlook communities lacking purchasing power or direct market influence (André et al., 2018; Radu, 2021; Cohen & Suzor, 2024; Ulnicane, 2024). The Right to AI maintains that these gaps warrant structured stakeholder participation to include marginalized voices and address power asymmetries.
@@ Appendix B Implementation Path
Urban planning regularly involves public input, though execution can vary (Jacobs, 1961; Sieber et al., 2024b). Projects like MID-Space used iterative community annotation to inform planning tools (Nayak et al., 2024), revealing how structured feedback loops might help integrate diverse local needs (Sloane et al., 2022).

### v2

@@ 1 Introduction
We use the term “Right to AI” to emphasize the collective governance dimension of AI oversight. This governance-oriented right is broader than more familiar claims such as the right to be forgotten, the right to explanation, or the right to contest AI decisions (Kaminski, 2019; Kaminski & Urban, 2021; Zhang et al., 2024). These latter rights, while important, speak mostly to individual entitlements to correct or clarify AI outputs. By contrast, the Right to AI we propose extends beyond mitigating harms to actively co-shaping the objectives, data practices, risk thresholds, and ethical principles of AI infrastructures. In this sense, it is more accurately viewed as a “power right” (Wenar, 2023) to guide AI’s development, rather than a narrower right to information or redress.
@@ 2.4 Grassroots Engagement
Previous attempts to incorporate participation in AI governance include user feedback loops in recommender systems, collaborative training data annotation, and community reviews of AI outputs (Gerdes, 2022; Zaidan & Ibrahim, 2024; Huang et al., 2024a). Innovative proposals like “jury-based” or “constitutional” approaches also engage diverse groups in AI ethics and policy deliberations (Gordon et al., 2022; Bai et al., 2022b; Sorensen et al., 2024). However, these methods are nascent and face scalability, political, market, resource, and institutional challenges (Saheb & Saheb, 2024; Zaidan & Ibrahim, 2024; Huang et al., 2024a).
Yet participation alone is not inherently empowering. It can become tokenistic—what some call “participation-washing”—when stakeholders are invited without real decision-making power or follow-through. In high-stakes contexts, such as healthcare or criminal justice, participation may also be constrained by the need for technical oversight or legal accountability. These realities suggest that participatory frameworks must be adapted to the specific risks, knowledge demands, and institutional capacities of each domain. Genuine inclusion requires not just the presence of diverse voices, but mechanisms that translate deliberation into influence.
@@ 3 Arguments for AI as Societal Infrastructure
A central premise of the Right to AI is that AI increasingly functions as societal infrastructure, comparable to utilities or educational systems. Viewing AI as societal infrastructure aligns with established frameworks of public goods, commons governance, and socio-technical systems (Ostrom, 1990; Graham & Marvin, 2001). Infrastructure commonly exhibits three properties: (i) broad societal impact, (ii) an essential role in daily life, and (iii) a requirement for collective management (North, 1990; Davern et al., 2017). AI, particularly foundation models shaping decisions having to do with employment, credit scoring, and public discourse, for example, arguably meets these criteria.
@@
Technologies that mediate access to financial systems, public services, and employment increasingly function as core societal infrastructure (North, 1990). AI-driven decision-making in credit approval, job screening, and social welfare administration underscores its role in structuring life chances (Eubanks, 2018; Benjamin, 2019). The opacity of these systems necessitates governance mechanisms akin to those regulating financial and legal infrastructures (Ulnicane, 2024).
@@
AI-based systems shape social interactions, political communication, and institutional trust (Gillespie, 2018; Kitchin, 2023). Like other infrastructures, AI is not neutral; it embeds political, economic, and cultural assumptions that shape its societal consequences (Crawford, 2021; Angwin et al., 2022). Without participatory oversight, AI risks reinforcing inequities rather than serving as a mechanism for collective well-being.
@@ 4.2 Social Justice and Pluralism
Machine learning models typically generalize from large datasets, which may fail to capture minority values or nuanced cultural norms (Goodfellow et al., 2016; Gebru et al., 2021; Dastin, 2022). As a result, marginalized voices risk erasure or misrepresentation (Raz, 1999; Beer, 2016; Bondi et al., 2021; Zhang et al., 2025). The Right to AI entails inclusive governance structures that protect pluralism by ensuring that multiple moral and cultural frameworks inform system design (Lefebvre, 1968; Costanza-Chock, 2020). This pluralistic perspective challenges any hegemonic assumption that there is a single “correct” data-driven solution (Fraser, 1995; Kitchin, 2023; Sorensen et al., 2024).
@@ 5.1 Consumer-Based (Minimal Right to AI)
5.2 Private Organization-Led
@@
5.3 Government-Controlled
@@ 5.4 Citizen-Controlled (Maximal Right to AI)
Citizen-controlled governance envisions significant communal authority over AI systems; however, this does not imply the exclusion of domain experts or the adoption of a uniform approach in every context. In critical fields such as healthcare and aviation, for instance, broad participation must be balanced with deep technical expertise to maintain safety and reliability. In these high-stakes domains, effective citizen control may take a hybrid form, wherein communities shape overall values and objectives while specialists guide specific technical parameters. Thus, citizen sovereignty in AI does not preclude expert collaboration; rather, it enables stakeholders to determine when and how specialized knowledge interacts with collective oversight.
Figure 2 illustrates how agency, transparency, inclusivity, and governance structures vary across these tiers. This progression also highlights the transition from instrumental consumerism to communal sovereignty, guiding evaluations of existing approaches and helping chart paths toward more participatory paradigms.
Figure 2: Progression in stakeholder power from minimal engagement (Consumer-Based) to robust self-governance (Communal Sovereignty). This categorization helps assess current initiatives and guide transitions toward more participatory models.
While our four-tier model is inspired by Arnstein’s seminal ladder of participation, it introduces AI-specific mechanisms for transitioning between tiers. For instance, the shift from Consumer-Based to Private Organization-Led may occur when communities adopt structured feedback channels and partner with companies to revise product roadmaps or data-sharing agreements. A subsequent transition to the Government-Controlled tier might involve formal policy mandates requiring community audits or the creation of statutory councils with partial decision-making authority over AI deployments. Finally, the Citizen-Controlled tier relies on advanced institutional support—such as local or cooperative data trusts with legally enforceable ownership structures and educational programs that foster the technical competence necessary for meaningful oversight.
At each transition point, stakeholder roles evolve—from reacting to system outputs, to co-managing data and design objectives, to exercising decisive authority over governance processes. By outlining these transitions, we aim to illustrate how communities can incrementally acquire the capacity to shape AI systems, moving beyond surface-level consultation or feedback.
@@ 6 Lessons from Participatory Practices
Participatory AI initiatives in education, healthcare, urban planning, and software development often involve stakeholders with direct interests in AI-driven decisions (Jacobs, 1961; Lee et al., 2019; Zicari et al., 2021; Zhang & Aslan, 2021; Sieber et al., 2024b). For example, Co-Design of Trustworthy AI in Healthcare (Zicari et al., 2021) included patients, clinicians, and ethicists to expose biases in diagnostic tools, leading to enhanced accountability despite resource and expertise challenges. Urban planning efforts, such as MID-Space (Nayak et al., 2024), relied on iterative community annotation and mediation to address conflicting priorities (Sloane et al., 2022). These examples highlight both the promise of stakeholder inclusion and the structural, institutional, and practical barriers that may limit its impact.
@@
Across the nine case studies, success factors consistently included the early and sustained involvement of local stakeholders (e.g., language experts, community volunteers), transparent articulation of goals and benefits, and explicit acknowledgment of resource inequalities. For example, in the WeBuildAI case study (Lee et al., 2019), the framework enabled diverse participants—including donors, volunteers, recipient organizations, and staff—to collaboratively design a matching algorithm for donation allocation. Evaluation using historical data revealed that the algorithm produced more equitable outcomes than traditional human-led methods. Specifically, it reallocated donations in ways that better prioritized organizations serving communities with higher poverty rates, lower incomes, and limited food access. Participants reported increased trust in, and clearer understanding of, the algorithmic decision-making process.
@@ 7 Recommendations
Mobilize Researchers for Community Engagement
As machine learning researchers are well equipped to raise awareness and morally support their surrounding communities through communication and dialogue about AI, this responsibility translates into practical steps. Implementing the Right to AI would prompt researchers to engage more systematically with non-technical stakeholders. This could involve structuring datasets with transparent documentation, designing interfaces for community feedback, and integrating diverse perspectives into model objectives. Such a shift may introduce time and resource overheads. However, it also alters the dynamics of accountability and offers opportunities to mitigate biases and strengthen public trust. Balancing technical efficiency with meaningful public engagement may require interdisciplinary collaboration, new skill sets—such as facilitation—and iterative design cycles. Although demanding, this participatory approach can enhance the relevance and robustness of AI systems while reinforcing public confidence in their development.
Bridge Technical Gaps
Beyond political and normative considerations, substantial technical gaps remain in realizing the Right to AI across all tiers. Participatory systems must offer accessible, language-inclusive interfaces that accommodate diverse forms of engagement and varying abilities, without oversimplifying complex modeling decisions. Moreover, conflict-resolution protocols require both computational and sociotechnical research to systematically integrate diverse perspectives into model design—shifting from disaggregated to aggregated viewpoints in a transparent and traceable manner. The development of explainability and interpretability tools tailored to non-experts remains in its early stages. Finally, reliable methods are needed to validate the quality and relevance of community-contributed data, particularly in regions with limited technical capacity. Addressing these challenges is essential to ensure that the Right to AI becomes not merely aspirational, but operationalized in a sustainable and equitable manner.
@@ 8 Alternative Views
Some scholars and practitioners question whether broad-based participatory approaches to AI are feasible or desirable. A market-led perspective asserts that competition and consumer choice will naturally drive responsible AI (Dignam, 2020; de Marcellis-Warin et al., 2022; Hadfield & Clark, 2023; Judge et al., 2024), though such models can overlook communities lacking purchasing power or market influence (André et al., 2018; Radu, 2021; Cohen & Suzor, 2024; Ulnicane, 2024). The Right to AI maintains that these gaps warrant structured stakeholder participation to include marginalized voices and address power asymmetries.
@@ 9 Conclusion
Consolidated Overview
Contemporary discourse on AI governance encompasses a broad spectrum of policy proposals, ethical guidelines, and technical approaches aimed at aligning AI systems with societal values (Mishra, 2023; Zaidan & Ibrahim, 2024; Sorensen et al., 2024). Major institutions such as the OECD and the European Union have introduced frameworks for responsible AI development, often emphasizing fairness, accountability, and transparency (Saheb & Saheb, 2024; Zhang et al., 2025). However, these initiatives vary widely in design and enforceability across different regions.
For instance, while the European Union’s proposed AI Act imposes legally binding obligations for high-risk AI systems (European Union, 2024), other jurisdictions frequently rely on voluntary guidelines or market-driven standards (Kerry et al., 2025). In the United States, the now-rescinded White House Blueprint for an AI Bill of Rights was primarily composed of non-binding principles, and no comprehensive replacement has emerged (The White House, 2025). In Asia, countries such as Singapore emphasize industry consultation and codes of practice, whereas Japan’s guidelines seek to harmonize technological innovation with broader societal goals (Kerry et al., 2025).
Several developing countries have drafted national AI strategies or joined initiatives like the African Union’s Continental Strategy for AI, but often face structural and financial barriers that constrain robust public participation (Bernstein & Bekheit, 2024). Overall, these disparate regulatory models illustrate an ongoing tension between top-down or expert-led approaches and a growing demand for more inclusive governance. The Right to AI advanced in this paper seeks to bridge these gaps by advocating participatory frameworks, human-centered design, and collective data ownership, thus complementing rather than merely supplanting existing governance structures across diverse legal and cultural contexts.
This paper’s core argument can be summarized in three steps: first, that AI increasingly functions as a form of societal infrastructure; second, that individuals and communities hold a right to help shape and govern the infrastructure that affects their lives (Benkler, 2006); and third, that therefore a “Right to AI” naturally follows. As elaborated in the main text, Section 3 advances the conceptual framing of AI as infrastructure, while Section 4 outlines four normative grounds—ethical, political, epistemic, and institutional—that justify this right.
In brief, we argue that:
  1. 1.
    to protect democratic legitimacy, impacted communities must have meaningful decision-making power over AI design;
  2. 2.
    social justice demands attention to marginalized voices;
  3. 3.
    epistemic autonomy requires safeguards against narrow or monolithic knowledge curation by AI;
  4. 4.
    data, which is essential for AI, is socially produced and thus merits communal oversight.
In grounding these claims, our framework views AI as not merely a product of private innovation but an infrastructure of public consequence, and it aligns with global calls for more inclusive and participatory technological governance (Crawford, 2021).
One might wonder if the Right to AI merely collapses into a broader right to infrastructure. While it does share foundational principles with other governance rights over public goods, AI’s particular characteristics—such as algorithmic opacity, global data reliance, and evolving automation capabilities—warrant a dedicated framework (Cohen, 2019). The Right to AI thus refines a general right to shape infrastructure, specifying the necessary mechanisms to address the unique ethical, technical, and political complexities posed by AI.
Some scholars and industry leaders dispute the premise that AI is or should be treated as public infrastructure, arguing that AI’s intangibility sets it apart from utilities like water or electricity. Others emphasize market-led solutions, suggesting that competition will naturally encourage responsible AI. However, the infrastructural viewpoint spotlights the breadth and depth of AI’s societal reach, from healthcare to education to political discourse (Plantin et al., 2018). This, we contend, justifies a governance paradigm akin to that of publicly regulated utilities. Treating AI as infrastructure thus reframes oversight as a collective undertaking, rather than a question of consumer choice or proprietary rights alone.
@@ Appendix B Implementation Path
Urban planning regularly involves public input, though execution can vary (Jacobs, 1961; Sieber et al., 2024b). Projects like MID-Space used iterative community annotation to inform planning tools (Nayak et al., 2024), revealing how structured feedback loops might help integrate diverse local needs (Mushkani et al., 2025).